# Value Alignment of Social Media Ranking Algorithms

### v1

@@ 6.2.1. What values do users select to re-rank their feeds?
Configured value weights reflect personal values. One hypothesis is that participants adjust sliders to reflect their personal value priorities. To test this, we compute the Pearson correlation ($r$) between each participant’s personal value rankings, measured via the Portrait Value Questionnaire, and their selected slider weights. The results support this hypothesis: $93.6\%$ of participants’ selected sliders are positively correlated with their personal values, with a mean correlation of $r=0.39\pm 0.25$. While only $38.3\%$ of participants show individually significant correlations ($p<0.05$), a one-sample $t$-test on Fisher’s $z$-transformed coefficients confirms that, overall, the positive alignment is highly significant (p<0.001p<0.001). [^9] Since each participant’s correlation is based on at most 19 values, individual tests are underpowered. Thus, many positive correlations do not reach the significance threshold despite being moderate in size. During our think-aloud studies, we observe participants referencing their personal values when selecting sliders. For instance, P2 chose to uprank “Tradition”, explaining, “Tradition is something that I have been brought up with … so it’s occasionally nice to see people talking about their homeland [on Twitter/X].
@@
6.2.2. How do value-aligned feeds compare to engagement feeds?
@@
Finally, we analyze what values are being promoted versus deprioritized in the re-ranked feeds. Using the method from Sec. 6.2.2, we compute the change in value strength from the participant’s engagement feed to the value-ranked feed for each testing round (Fig. 8). Here, the maximum possible change in value strength is $42.4$, which is equivalent to going from an engagement feed where the value is not expressed to a re-ranked feed where each post strongly expresses the value. Overall, value strength increases across all dimensions as more value-laden content is promoted during re-ranking. There is a larger boost for values expressing self-transcendence ($4.47\pm 1.053$) as opposed to self-enhancement ($1.27\pm 1.13$) and openness to change ($3.83\pm 2.33$) over conservation ($1.31\pm 1.13$). Value-aligned feeds contained more content concerning the welfare or interests of others rather than posts emphasizing the pursuit of individuals’ interests.
6.2.3. When do users want value based controls?
Finally, we examine when and why users want to use value-based controls.
[…]
@@ 7. Discussion
Our work envisions a social media ecosystem where basic human values, with their inherent trade-offs, are directly incorporated into feed ranking algorithms. We show that ranking feeds by values can indeed be done, and we present an approach for doing so that can be adopted by industry practitioners. The method we propose consists of multiple steps: identifying values in content, scaling up the coding process using large language models, eliciting value priorities from users, and incorporating these priorities when ranking feeds. Our evaluation confirms that users can distinguish between value-aligned feeds ranked using this approach and ones that are not value-aligned.
Our design decisions and findings carry implications for future technical endeavors that aim to adopt or adapt our approach for ranking feeds by values. Here we outline key considerations for future work including potential challenges, opportunities for improvement, and ethical considerations.
7.1. Value Bubbles
Empowering users to curate feeds that align with their values raises the consideration of how and whether this agency could trap them in “value bubbles”, limiting their exposure to content that reflects different values. It is conceivable that end-user value-based tools might make it easier for us to avoid seeing complementary points of view.
Simultaneously, we must consider: our feed algorithms already implicitly embed values (Bernstein et al., 2023), and we have no control over them. Has the present lack of agency and surrendering to a curation model that produces value-laden feeds as a second degree effect of optimizing for engagement metrics already led to a value monoculture—one where certain individualistic values are heightened for all users (Kleinberg and Raghavan, 2021)? If the current ecosystem is already limiting diversity in values, might giving users more control in fact offer a chance to reintroduce that diversity?
Value-sensitive content curation can pave the way for puncturing existing filter bubbles by identifying content that bridges divides. For instance, while users may resist content that is not aligned with their views on topics such as politics or otherwise, they could be more receptive if it is presented in a way that aligns with their core values (Törnberg et al., 2023).
Another argument for why it might be beneficial for users, as opposed to platforms or other centralized authorities, to have oversight on what values are reflected in their feeds is the inadequacy of centralized moderation to address the unique needs of every user or community (Jahanbakhsh et al., 2022a) and the risk to further marginalize those already overlooked. In addition, deferring the judgment to users treats them as autonomous individuals and upholds their moral rights to freedom of listening and against compelled listening (Corbin, 2009).
While it does not downplay the risk of value bubbles, it is worth noting that, years later, we still do not have strong empirical evidence that social media actually cause filter bubbles. For example, users of social media are exposed to more cross-cutting political news than those who do not use social media (Fletcher and Nielsen, 2018), and increasing exposure to cross-cutting views on social media had either no or very small impacts on affective polarization (Nyhan et al., 2023).
All of this given, it is still worth considering how we might mitigate the risk of enabling value bubbles. One approach would be to draw on social choice or other democratic methods to place bounds on how much users can value-shape their feeds. For example, an algorithm might enforce that people can only influence a value’s prevalence to within some percentile of the overall population distribution, making it difficult to completely mute competing values. Another approach is to implement bridging algorithms (Törnberg et al., 2023) that seek out content that is counter-valued but still broadly appreciated.
One final risk bears consideration and mitigation. What if a platform or state actor decides to utilize these techniques to enforce specific values on its populace? In these cases, it may still be useful to empower end users to rerank content privately on the client side (Kolluri et al., 2025). Finally, this method helps make it possible to audit the values getting amplified by a feed algorithm, so that these interventions can be detected.
7.2. Designing Interfaces for Eliciting Users’ Value Priorities
The interface we designed to elicit users’ value priorities used 19 sliders, one for each value. Yet this approach represents just one point in a larger design space. Exploring alternatives requires grappling with trade-offs between interface complexity, users’ self-efficacy or perceived control over their feed, and the recognizability of the resulting feeds. As surfaced in our think-aloud study, some participants expressed a desire for an interface that allowed them to describe their values in natural language. This is a promising direction, but it raises the question of how best to guide users’ reflection, as people may struggle to articulate their values when presented with a blank slate. Another insight from participants was that how much they cared about a given value was influenced by the anticipated presence of other values in the same content. This kind of conditional preference demands finer-grained control than what is currently afforded by the sliders which are adjusted independently of each other. To support this expressivity without overwhelming users, we might consider designing value clusters, i.e., groups of values that are often sought together, or even enabling users to construct their own clusters using logical operators (e.g., and, or).
7.3. Navigating Value Trade-offs
The system of values that we adopt—Schwartz’s theory of Basic Human Values—is a complete value system in the sense that it seeks to both articulate a full design space and also describe to what extent these values are conflicting or complementary. To reiterate, other value systems could be used—we make no claim that Schwartz’s theory is the only valid approach. We find Schwartz fit for our purpose because its values are few enough in number that they can be integrated into user interface controls, it is widely studied and replicated, and it provides the value similarities and differences inherent to a complete value system. Other value systems fitting these criteria could replace Schwartz in our approach. While we evaluated our approach in the context of Schwartz’s values, we believe that the stages that it comprises—from operationalizing values in tweets, to scaling value annotations, to ranking content based on values—can generalize to values beyond those in Schwartz’s system.
Schwartz’s value system introduces limitations, however. Not every value in Schwartz’s theory is immediately understandable to a lay user. This challenge led us to reword several of the value names in our interfaces in consultation with a cultural psychologist, to aid usability. In addition, while Schwartz’s Basic Human Values are intended to be comprehensive, there may be specific, nuanced, or culturally unique values that are difficult to fully capture within this framework, particularly in the context of social media. Therefore, there is an opportunity for extending our work to other values or value systems, and perhaps even those that individuals or communities have identified as important through their own experiences (Weld et al., 2024, 2022), rather than constraining them to a predefined top-down set of values. Future work should investigate offering users a compilation of these various values and interfaces for capturing them, allowing for a more tailored content curation.
@@ 7.3. Navigating Value Trade-offs
In Study 2, users often used our value elicitation interfaces to change the prominence of many values in their feed simultaneously. When most of the values are activated, it can result in the feed being influenced in different, and sometimes conflicting, directions. Our value elicitation interface in fact masked these conflicts by allowing a user to change their desired prominence of each value independently of other values. Nonetheless, participants could distinguish which feeds were aligned with their intentions even in the complex scenario when they had to reason over potentially conflicting values. This finding points to potential for developing tools that help users navigate and balance these conflicts. Future work should investigate how people reason about conflicts in value objectives, how we might better communicate these trade-offs to them, and how to elicit and help users come to a decision about their ideal value priorities. These questions are, in fact, larger than social media. With growing concerns about whose values large language models encode and output (Santurkar et al., 2023; Durmus et al., 2025), these questions are also highly relevant to the AI alignment community.
7.4. Balancing Engagement Signals with Values
@@
The profitability of engagement-based ranking for social media companies suggests that value-based ranking, if adopted, would be incorporated alongside engagement metrics in the content curation process. Early work offered promising results that ranking by pro-democratic attitudes does not negatively impact engagement (Jia et al., 2024). While our work applied value ranking at the last stage of the pipeline (i.e., on content that was already filtered for engagement), where in the pipeline values should be considered is a question that warrants investigation. The distribution of values across all the content that is posted on social media is likely uneven, and the distribution may further differ in posts from a user’s connections or posts that receive high engagement. If industry practitioners adopt our approach, they should investigate these variations in distribution of values and the impact of incorporating value ranking earlier in the pipeline, not only on the status-quo metrics of user engagement and retention, but also on longer term metrics such as user well-being and attitude. However, in practical terms, we acknowledge that platforms often allow engagement to dominate any ultimate decisionmaking on what gets launched (Christin et al., 2024).
7.5. Democratization of Value Classification
@@
While Schwartz’s theory provides a research-grounded vocabulary for discussing values, it does not fully resolve the subjectivity involved in interpreting them. Which specific values people associate with a piece of content — and to what extent the value is evoked for them — can vary due to the subjective nature of values. This variation in perception may be influenced by their background or the values they personally prioritize, among other factors (Haidt and Joseph, 2004). The labeling scheme in our approach involved providing the list of Basic Human Values, their definitions, and examples of tweets coded by the researchers, which were then used for in-context learning by a large language model. In general, our results suggest that this approach worked: participants recognized the value-aligned feeds that they created. However, differences between people in how we perceive value expressions in content might explain why in some cases our users did not recognize their value-aligned feeds. Future work should investigate the factors driving the difference in perceived values in content and build classifiers customized for different groups of users that capture their unique experiences and perceptions. This approach could ensure that value-aligned feeds are more relevant and resonant for various user clusters.
7.6. Limitations
@@ Acknowledgements.
We thank members of the Stanford HCI Group and Social Media AI Reading Group for their helpful feedback. This work was sponsored by the Hoffman-Yee Research Grants at the Stanford Institute for Human-Centered Artificial Intelligence (HAI). Dora Zhao is supported in part by the Brown Institute for Media Innovation. Tiziano Piccardi is supported by Swiss National Science Foundation (Grant P500PT-206953). Zachary Robertson is supported by a Stanford School of Engineering fellowship. Sanmi Koyejo acknowledges support from NSF 2046795 and 2205329, the MacArthur Foundation, Stanford HAI, and Google Inc.
@@ A.1. Method
Value Labeling Prompt Consider the following set of concepts, listed as !<CONCEPT>! : !<DEFINITION>!: \${conceptDefinitionsStr} For each concept, rate whether the tweet reflects the concept on a Likert scale with options [0, 1, 2, 3, 4, 5, 6], where: 0 = This post does not reflect this concept at all 1 = This post reflects this concept a little bit 6 = This post reflects this concept strongly A concept can be reflected if the tweet supports the concept. Output: One JSON dictionary of the following format. Rating is a dictionary of key value pairs, with each key being a concept and each value your rating for that concept: ”Rating”: ”Concept”: !¡RATING¿! Example 1: Tweet: I’ve always believed in the power of research to save lives and ensure Americans get the care they need. Starting today, the first-ever White House Initiative on Women’s Health Research will work towards that goal, changing how we approach and fund women’s health research. {”Rating”: {”Reputation”: 0, ”Power”: 0, ”Wealth”: 0, ”Achievement”: 0, ”Pleasure”: 0, ”Independent thoughts”: 0, ”Independent actions”: 0, ”Stimulation”: 0, ”Personal security”: 0, ”Societal security”: 0, ”Tradition”: 0, ”Lawfulness”: 0, ”Respect”: 0, ”Humility”: 0, ”Responsibility”: 5, ”Caring”: 4, ”Equality”: 5, ”Connection to nature”: 0, ”Tolerance”: 3}} Example 2: Tweet: This little kid was carrying a white flag, and now he’s dead. This guy was also carrying a white flag, and he’s been shot. I’m here filming for you, and I’m in a lot of danger as well. People holding white flags are trying to come out, and are scared from the snipers. If the claims of civilians with white flags getting INTENTIONALLY targeted is true, this would be a WAR CRIME. {”Rating”: {”Reputation”: 0, ”Power”: 0, ”Wealth”: 0, ”Achievement”: 0, ”Pleasure”: 0, ”Independent thoughts”: 0, ”Independent actions”: 0, ”Stimulation”: 0, ”Personal security”: 0, ”Societal security”: 0, ”Tradition”: 0, ”Lawfulness”: 0, ”Respect”: 0, ”Humility”: 0, ”Responsibility”: 0, ”Caring”: 6, ”Equality”: 5, ”Connection to nature”: 0, ”Tolerance”: 3}} Tweet:

### v2

@@ 6.1. What values do users select to re-rank their feeds?
Configured value weights reflect personal values. One hypothesis is that participants adjust sliders to reflect their personal value priorities. To test this, we compute the Pearson correlation ($r$) between each participant’s personal value rankings, measured via the Portrait Value Questionnaire (PVQ), and their selected slider weights. For each of the nineteen values, the PVQ returns a continuous score ranging from 5-5 to $5$, which represents the relative importance of the value to the individual. The results support our hypothesis that participants’ slider selections match their personal values: $93.6\%$ of participants’ selected sliders are positively correlated with their personal values, with a mean correlation of $r=0.39\pm 0.25$. While only $38.3\%$ of participants show individually significant correlations ($p<0.05$), a one-sample $t$-test on Fisher’s $z$-transformed coefficients confirms that, overall, the positive alignment is highly significant (p<0.001p<0.001). [^10] Since each participant’s correlation is based on at most 19 values, individual tests are underpowered. Thus, many positive correlations do not reach the significance threshold despite being moderate in size.
@@
Value weights mirror Schwartz’s trade-off structure. We select Schwartz’s value system as it presents a design space of values containing trade-offs, and enumerates values that are complementary versus those in tension in a structured fashion (i.e., based on distance in the circumplex). If users’ value selections reflect these theoretical relationships, we would expect that users are unlikely to simultaneously uprank values that are in tension or downrank values that are complementary. To examine how these trade-offs manifest in users’ configured value weights, we compute the Pearson correlation ($r$) between the distance separating two values in Schwartz’s circumplex and the product of the weights the user assigned to them. We define distance $d$ as $1-\frac{2\theta}{\theta_{\max}}$ where $\theta$ is the angular distance between two values in the circumplex (Fig. 2) and $\theta_{\max}$ is the maximum possible angular distance. This metric ranges from -1 for values at the opposite ends of the circumplex to 1. After applying a Fisher zz–transformation, we observe a significant negative relationship between our distance metric and value weights with a mean correlation coefficient of 0.15-0.15 ($p<0.001$). This result indicates that users tend to assign higher weights to complementary values (e.g., “Caring” and “Universal concern”) and lower weights to values in tension (e.g., “Tradition” and “Self-directed thought”), consistent with engaging in meaningful value trade-off reasoning.
6.2. How do value-aligned feeds compare to engagement feeds?
@@
Bar chart showing the average strength of each value in participants’ engagement-aligned feeds. Bar chart showing the average strength of each value in participants’ engagement-aligned feeds. Values with a personal focus, such as ”Stimulation”, ”Achievement”, ”Hedonism”, appear more frequently than social-focus values, like ”Caring” or ”Humility.”
@@
Bar chart showing the mean change in value strength from participants’ engagement-aligned feeds to the value-aligned feed. Bar chart showing the mean change in value strength from participants’ engagement-aligned feeds to the value-aligned feed. There is a greater increase for values related to Self-Transcendence (e.g., ”Caring”, ”Universal Concern”) and Openness to Change (”Self-directed actions”, ”Self-directed thoughts”).
Finally, we analyze what values are being promoted versus deprioritized in the re-ranked feeds. Using the method from Sec. 6.2, we compute the change in value strength from the participant’s engagement feed to the value-ranked feed for each testing round (Fig. 8). Here, the maximum possible change in value strength is $42.4$, which is equivalent to going from an engagement feed where the value is not expressed to a re-ranked feed where each post strongly expresses the value. Overall, value strength increases across all dimensions as more value-laden content is promoted during re-ranking. There is a larger boost for values expressing self-transcendence ($4.47\pm 1.053$) as opposed to self-enhancement ($1.27\pm 1.13$) and openness to change ($3.83\pm 2.33$) over conservation ($1.31\pm 1.13$). Value-aligned feeds contained more content concerning the welfare or interests of others rather than posts emphasizing the pursuit of individuals’ interests.
@@ 7. Discussion
Our work envisions a social media ecosystem where basic human values, with their inherent trade-offs, are directly incorporated into feed ranking algorithms. In this section, we highlight both the promise and the complexity of building value-aligned social media feeds, drawing on the design decisions we worked through. Designing real systems that support coherent value alignment requires design across the entire pipeline, from how to design interfaces that enable users to express their values (Sec. 7.1), to how values could be incorporated into ranking infrastructure (Sec. 7.2), to how the system interprets and classifies values in content (Sec. 7.3). Below we outline several design considerations for building such systems in practice.
7.1. Eliciting Value Priorities
The interface we designed to elicit users’ value priorities used 19 sliders, one for each value. Yet this approach represents just one point in a larger design space. Exploring alternatives requires grappling with trade-offs between interface complexity, users’ perceived control over their feed, and the recognizability of the resulting feeds. One promising direction is eliciting values via natural language, which provides users with greater expressivity. For example, users might want for values to be activated conditionally or in combination with each other, such as wanting to see content that expresses only societal security and caring in political posts. This level of expressivity is not possible with our current interface but could be elicited via natural language. An open question for natural language interfaces is how to best guide users’ reflection, such as through free-form text entry or a multi-turn interview setting (Li et al., 2025).
@@ 7.1. Eliciting Value Priorities
A broader challenge in interactive value elicitation is that users must anticipate how their selections will translate into ranking behavior. This task becomes even more challenging when users must reason over how multiple values interact with each other, as the feed may be influenced in different, and sometimes, conflicting directions. Our value elicitation interface in fact masked some of these conflicts by allowing users to adjust each value independently. Nonetheless, participants could still distinguish which feeds were aligned with their intentions even when they had to reason over potentially competing values. This finding points to potential for developing tools that help users navigate and balance these conflicts. Future work should investigate how people reason about conflicts in value objectives, how we might better communicate these trade-offs, and how to support users in deciding on their ideal value priorities. Furthermore, designing interfaces that clarify input-output mappings — e.g., through examples — may enable more meaningful control. One promising direction is to support structured forms of expressivity, such as value clusters (groups of values often sought together) or user-defined combinations using logical operators (e.g., and, or), which can provide finer-grained control and make users’ intentions more explicit.
Finally, while Schwartz’s theory treats values as relatively stable principles, he also reports that people’s values might shift over time in response to life events, changing social contexts, or evolving media habits. For this reason, value-aligned feed interfaces should support revisiting and adjusting value settings when users choose to do so, much like existing controls for interests or personalization. This explicit control maintains user autonomy without requiring continuous monitoring or inference of values. At the same time, platforms already support preference refinement through lightweight, optional feedback mechanisms (e.g., “show me more of this” and “I liked this”). Value-aligned feeds could incorporate analogous interactions, such as letting users indicate when a post “aligns with my values”, which could be a low-effort way for users to refine their values over time without rearticulating their entire profile.
7.2. Incorporating Values into Feed Ranking Algorithms
@@
The profitability of engagement-based ranking for social media companies suggests that value-based ranking, if adopted, would be incorporated alongside engagement metrics in the content curation process. Our work applied value-based reranking at the last stage (i.e., on content that was already filtered for engagement), but real-world systems may incorporate values at earlier stages of candidate generation and relevance estimation. Because the distribution of values in posted content is uneven and further filtered by connection networks and engagement dynamics, introducing value-based ranking earlier in the pipeline can more powerfully reshape what becomes visible. If industry practitioners adopt our approach, they should evaluate the impacts of different insertion points on users, not only on the status quo metrics of user engagement and retention, but also on longer term metrics such as user well-being and attitude. However, in practical terms, we acknowledge that platforms often allow engagement to dominate any ultimate decision-making on what gets launched  (Christin et al., 2024).
Another consideration is how heavily end-user values ought to shape the feed — a question that connects to broader debates about filter bubbles on social media that arise through the dynamics of selective exposure and personalization (Bozdag and Van Den Hoven, 2015; Kitchens et al., 2020). Users may end up only seeing posts that reinforce their views, limiting their exposure to other values. One way to mitigate this risk is to draw on social choice or other democratic methods to place bounds on how much users can value-shape their feeds. For example, an algorithm might enforce that people can only influence a value’s prevalence to within some percentile of the overall population distribution, making it difficult to completely mute competing values. Another approach is to implement bridging algorithms (Törnberg et al., 2023) that seek out content that is counter-valued but still broadly appreciated.
A third consideration concerns the ranking function itself. In our implementation, value-based ranking relied on linear weights over content-user value scores. While this choice provides transparency and aligns with standard industry practice of using linear objective functions in ranking, it constrains how complex interactions among values can affect ranking outcomes. More expressive approaches, such as non-linear aggregation or contextualized decision rules, may capture richer value interactions, though they introduce interpretability and governance challenges that merit further study.
7.3. Choosing Value Systems to Model
@@
While Schwartz’s theory provides a research-grounded vocabulary for discussing values, individuals differ in which values a piece of content evokes for them, in part due to their backgrounds, experiences, or values that they personally prioritize among other factors (Haidt and Joseph, 2004). The labeling scheme in our approach involved providing the list of Basic Human Values, their definitions, and examples of tweets coded by the researchers, which were then used for in-context learning by a large language model. In general, our results suggest that this approach worked: participants recognized the value-aligned feeds that they created. However, variation in perceived value expressions likely contributed to the instances in which users did not recognize their value-aligned feeds. Future work should investigate the factors driving the difference in perceived values in content and build classifiers customized for different groups of users that capture their unique experiences and perceptions. This approach could ensure that value-aligned feeds are more relevant and resonant for various user clusters.
These perceptual differences also raise a broader question of which value system should a feed use as its underlying representational space. Although we treat Schwartz’s theory of Basic Human Values as a practical foundation, it is not the only valid value system. We adopt Schwartz because it is a complete value system, in the sense that it seeks to both articulate a full design space and also describe to what extent these values are conflicting or complementary. Additionally, its values are few enough in number that they can be integrated into user interface controls, it is widely studied and replicated, and it provides the value similarities and differences inherent to a complete value system. Other value systems fitting these criteria could replace Schwartz in our approach. While we evaluated our approach in the context of Schwartz’s values, we believe that the stages that it comprises — from operationalizing values in tweets, to scaling value annotations, to ranking content based on values — can generalize to values beyond those in Schwartz’s system.
Nonetheless, relying on a predefined value system introduces practical limitations. Not every value in Schwartz’s theory is immediately understandable to a lay user. This challenge led us to reword several of the value names in our interfaces in consultation with a cultural psychologist, to aid usability. In addition, even a comprehensive value system may fail to capture values that are culturally specific, nuanced, or especially salient in the context of social media. Therefore, there is an opportunity for extending our work to other values or value systems, and perhaps even those that individuals or communities have identified as important through their own experiences (Weld et al., 2024, 2022), rather than constraining them to a predefined top-down set of values. Future work should investigate offering users a compilation of these various values and interfaces for capturing them, allowing for a more tailored content curation. These questions are, in fact, larger than social media. With growing concerns about whose values large language models encode and output (Santurkar et al., 2023; Durmus et al., 2025), these questions are also highly relevant to the AI alignment community.
8. Limitations
@@
In our work, we use recognizability of the resulting feed as the primary evaluation metric for value alignment. We acknowledge that recognizability does not capture all possible forms of alignment inside the model. However, it is the most direct and interpretable measure of whether the system’s input–output mapping remains meaningful from the user’s perspective.
@@ Acknowledgements.
We thank members of the Stanford HCI Group and Social Media AI Reading Group for their helpful feedback. This work was sponsored by the Hoffman-Yee Research Grants at the Stanford Institute for Human-Centered Artificial Intelligence (HAI) and NSF 2403435. Dora Zhao is supported in part by the Brown Institute for Media Innovation and the Paul & Daisy Soros Fellowship for New Americans. Tiziano Piccardi is supported by Swiss National Science Foundation (Grant P500PT-206953). Zachary Robertson is supported by a Stanford School of Engineering fellowship. Sanmi Koyejo acknowledges support from NSF 2046795 and 2205329, the MacArthur Foundation, Stanford HAI, and Google Inc.
@@ A.1. Method
Consider the following set of concepts, listed as !<CONCEPT>! : !<DEFINITION>!:
\${conceptDefinitionsStr}
For each concept, rate whether the tweet reflects the concept on a Likert scale with options [0, 1, 2, 3, 4, 5, 6], where:
  • 0 = This post does not reflect this concept at all
  • 1 = This post reflects this concept a little bit
  • 6 = This post reflects this concept strongly
A concept can be reflected if the tweet supports the concept.
Output: One JSON dictionary of the following format. Rating is a dictionary of key value pairs, with each key being a concept and each value your rating for that concept:
”Rating”: ”Concept”: !¡RATING¿!
Example 1:
Tweet: I’ve always believed in the power of research to save lives and ensure Americans get the care they need. Starting today, the first-ever White House Initiative on Women’s Health Research will work towards that goal, changing how we approach and fund women’s health research.
{”Rating”: {”Reputation”: 0, ”Power”: 0, ”Wealth”: 0, ”Achievement”: 0, ”Pleasure”: 0, ”Independent thoughts”: 0, ”Independent actions”: 0, ”Stimulation”: 0, ”Personal security”: 0, ”Societal security”: 0, ”Tradition”: 0, ”Lawfulness”: 0, ”Respect”: 0, ”Humility”: 0, ”Responsibility”: 5, ”Caring”: 4, ”Equality”: 5, ”Connection to nature”: 0, ”Tolerance”: 3}}
Example 2:
Tweet: This little kid was carrying a white flag, and now he’s dead. This guy was also carrying a white flag, and he’s been shot. I’m here filming for you, and I’m in a lot of danger as well. People holding white flags are trying to come out, and are scared from the snipers. If the claims of civilians with white flags getting INTENTIONALLY targeted is true, this would be a WAR CRIME.
{”Rating”: {”Reputation”: 0, ”Power”: 0, ”Wealth”: 0, ”Achievement”: 0, ”Pleasure”: 0, ”Independent thoughts”: 0, ”Independent actions”: 0, ”Stimulation”: 0, ”Personal security”: 0, ”Societal security”: 0, ”Tradition”: 0, ”Lawfulness”: 0, ”Respect”: 0, ”Humility”: 0, ”Responsibility”: 0, ”Caring”: 6, ”Equality”: 5, ”Connection to nature”: 0, ”Tolerance”: 3}}
Tweet:
@@
Distribution of value labels across tweets The figure shows histograms for all nineteen values, indicating how frequently each value appears across 212,663 tweets. Most tweets contain no explicit value labels, with counts heavily concentrated at zero across all categories, demonstrating that values are sparsely expressed. Among the labeled values, Stimulation and Hedonism appear most frequently relative to others.